OpenStudy (anonymous):

a car of mass 800 kg is moving along a straight road under the action of constant resultant force of 1000N. find its speed when it has moved 40m from rest

OpenStudy (anonymous):

work=F*d F*d=1/2*m*v^2 1000*40=0.5(800)*v^2 v=10m/s